# Quantum correlations under decoherence

### v1

@@ I Introduction
Quantum states are fundamentally different from classical states. Any measurements made on one part of a bipartite state will make the results uncertain. The integral nature of quantum states is the reason for this randomness, not the measuring device. Only states which can be described by a classical probability distribution, are invariant under such local measurements. Perhaps the most striking feature of quantum mechanics is the quantum entanglement (QE). It is at the heart of quantum information processing, being a resource for doing tasks like quantum computing, quantum teleportation and quantum key distribution NiCh-book-2000 . Nonetheless, in the early age of quantum information science, some researchers questioned entanglement as the source of power for quantum computation MiLaSaKn-PRA-2002 . QE is not the only correlation which exists between different parts of a multi-party quantum system. Although pure states can be either uncorrelated or just entangled AdBrCi-JPA-2016 , it has been proven that certain quantum separable mixed states exhibit quantum correlations, the so-called quantum discord OlZu-PRL-2001 ; HeVe-JPA-2001 . In addition to the significant role of quantum correlations in efficient quantum communication and computational tasks, they are also important in figuring out cooperative phenomena in multi-party quantum systems. Among various beyond entanglement quantum correlations have already been identified, one can mention local quantum uncertainty (LQU) GiTuAd-PRL-2013 , measurement-induced nonlocality (MIN) and correlated quantum coherence LuFu-PRL-2011 ; HuFa-NJP-2015 . While the former is introduced as the minimum skew information WiYa-PNAS-1963 achievable on a single local measurement, the latter is interpreted as the maximum global effect caused by locally invariant measurements.
@@
Beside quantum correlations, quantum coherence BaPl-PRL-2014 , being cornerstone of quantum computation and quantum communication theory, has a fundamental and practical significance for developing the quantum mechanics Huetal-PR-2018 . Contrary to quantum correlations, being defined for systems with at least two parties, quantum coherence measures can also be defined for single systems. However, both arise from quantum superposition principle, thus being interrelated. In this regard, the link between quantum coherence and entanglement have already been identified StSiDhBeAd-PRL-2015 ; QiGaYa-JPA-2017 . Furthermore, the concept of “correlated coherence”(CC) Taetal-PRA-2016 has been introduced and its connection to quantum discord and entanglement has been argued. The idea is splitting the coherence contained in the global state into two portions, those stored locally in the marginal (reduced) states and that stored non-locally only in the correlated state, and then subtracting local portions from the correlated one. For X states, we’ll find that MIN equals CC.
@@
Since a quantum system cannot be completely isolated from its environment, its influence on the system’s dynamics must be taken into account. Two approaches have been developed to study dynamics of open quantum systems: the quantum operations formalism and the approach of master equations which each one has its own place NiCh-book-2000 . The latter strategy employs a differential equation, which in addition to the conventional unitary term for an isolated system, includes a non-unitary term responsible for interaction with the surrounding environment. In the decoherence process, one of the most striking results is the complete loss of entanglement in a finite time, the so-called “entanglement sudden death” (ESD) implying complete termination of entanglement after a finite time interval without a long-lasting smooth decrease YuEb-PRL-2004 ; YuEb-PRL-2006 ; EbYu-Sci-2007 . Mathematically, ESD is based on the cutoff in the definition of concurrence, but it is difficult to find a physical interpretation. However, its connection to energy transfer between the system and its surroundings has been discussed CuLiYi-PLA-2007 . On contrary, the environment-induced FiTa-PRA-2006 or decoherence-induced entanglement DaLi-JPB-2007 has been worked out. Here, environment plays a constructive role, in contrast to its customary destructive role, inducing QE. Furthermore, entanglement re-birth or revival has also been reported.
@@
In this work, we examine a system of two localized coupled qubits that that are in interaction with an external magnetic field. The interaction with the environment is introduced by a non-unitary term in the Lindblad equation. We are going to solve this equation analytically first for a number of initial states in the form X and then for a general X-shaped state; and study the evolution of different quantum correlations for different emission rates as well as different interaction intensities. It is imperative to determine whether quantum correlations beyond entanglement experience sudden death or not. Furthermore, it is important to conduct an investigation into the phenomenon of correlation induced by the environment. After all this, it will be time to consider the effect of temperature on quantum correlations. How does this parameter affect ESD?
@@ II Some Quantum Correlations; entanglement and beyond
In this section we briefly review QE, LQU and MIN. At the end, we consider CC and relate it to MIN. Different measures have been introduced for quantification of QE. One of the most famous measures of entanglement is the concurrence HiWo-PRL-1997 ; Wo-PRL-1998 which for a two-qubit state $\rho$ is defined as
@@
$\lambda_{i}$s being the eigenvalues of the Hermitian matrix $\sqrt{\sqrt{\rho}\tilde{\rho}\sqrt{\rho}}$ in non-increasing order. Here, $\tilde{\rho}$ is the the spin-flipped Hermitian density matrix given by $(\sigma_{y}\otimes\sigma_{y})\rho^{*}(\sigma_{y}\otimes\sigma_{y})$; $\sigma_{y}$ being the second Pauli matrix and $\rho^{*}$ the complex conjugate of $\rho$ which due to the Hermiticity of $\rho$ equals its transpose, $\rho^{*}=\rho^{T}$. Two matrices $S=\sqrt{\rho}\tilde{\rho}\sqrt{\rho}$ and $R=\rho\tilde{\rho}$ are connected through the invertible matrix $\sqrt{\rho}$ i.e., $R=(\sqrt{\rho})S(\sqrt{\rho})^{-1}$. Thus, they are similar, having the same spectrum. In this way, one sees $\lambda_{i}$s are the square roots of the eigenvalues of $S$ and thus the non-Hermitian matrix $R$. Note that since $\sigma_{y}\otimes\sigma_{y}$ is its own inverse, $(\sigma_{y}\otimes\sigma_{y})^{-1}=\sigma_{y}\otimes\sigma_{y}$ then $\tilde{\rho}$ is similar to $\rho^{*}$ which its eigenvalues are just those of $\rho$. Therefore, both $\rho^{*}$ and its similar matrix $\tilde{\rho}$ are non-negative. Now, since both $\rho$ and $\tilde{\rho}$ are non-negative, eigenvalues of the their product $R=\rho\tilde{\rho}$ are non-negative Bh-book-1997  [^1] See page 256.. Concurrence, defined in this way, becomes zero for separated states but takes the value one for maximally entangled states in $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ systems Dhetal-book-2017 . It has been argued that the concurrence is related to the entanglement of formation which itself is an entropic entanglement monotone HiWo-PRL-1997 ; Wo-PRL-1998 . In the basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$, a X-state has the form
@@
coming from properties of the density operator i.e., unit trace, Hermiticity and positivity. The concurrence (1) for the X-state (6) is then Waetal-JPB-2006
$\mathcal{C}$ $=$ $2\max\bigg{\{}0,|\rho_{14}|-\sqrt{\rho_{22}\rho_{33}},|\rho_{32}|-\sqrt{\rho_{11}\rho_{44}}\bigg{\}}.$ (8)
As this equation show, if and only if either $\rho_{22}\rho_{33}<|\rho_{14}|^{2}$ or $\rho_{11}\rho_{44}<|\rho_{23}|^{2}$ then the state (6) is entangled. Both conditions cannot be fulfilled simultaneously AlRaAl-PRA-2010 .
As another natural and widely used entanglement measure, negativity for a bipartite state $\rho$ is defined through the trace norm of the partially transposed density matrix ViWe-PRA-2002 ,
@@
where $\{|i\rangle\}$ and $\{|k\rangle\}$ are given local orthonormal basis for subsystems $A$ and $B$ respectively. It is noteable that the spectrum of the partial transposition of the density matrix is independent both of the choice of the local basis and also whether the partial transposition is performed over the party $A$ or $B$ PlVi-bookchap-2014 . One can easily check that $\rho^{T_{B}}$ is Hermitian
@@
where in the third equality we have used Hermiticity of the density operator $\rho$. Now from Hermiticity of $\rho^{T_{B}}$ and the fact that trace is independent of basis, one has that
@@
where in the third equality we have used the fact that for entangled two qubit states the partiality transposed state $\rho^{T_{B}}$ has only one negative eigenvalue MiGr-JPB-2004 which has been denoted by $\lambda_{\min}$ in Eq. (12) as it is the minimum eigenvalue. Some authors have defined negativity of two-qubit states as
@@
which for maximally entangled states where the spectrum of $\lambda$ is $\{-\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2}\}$, takes the largest value 1 ZyHoSaLe-PRA-1998 ; VeAuDeMo-JPA-2001 . For a mixed state of two qubits, the subject of our study, it has been proved that the negativity $\mathcal{N}^{\prime}$ of a state can never exceed its concurrence and is always greater than $\sqrt{(1-\mathcal{C})^{2}+\mathcal{C}^{2}}-(1-\mathcal{C})$; $\mathcal{C}$ being the concurrence of the state VeAuDeMo-JPA-2001 . Now, from Eqs. (12) and (13) we have that $\mathcal{N}=\mathcal{N}^{\prime}/2$ and thus from the results of VeAuDeMo-JPA-2001 we have that
@@
The logarithmic negativity (log-negativity) as an appropriate quantifier of QE in a bipartite quantum system is defined as ViWe-PRA-2002 ; Pl-PRL-2005
@@
Now we consider quantum correlations beyond QE. Based of the notion of skew information, authors of GiTuAd-PRL-2013 have proposed a discord-like scheme for characterizing non-classical correlation via LQU. In this way, LQU has been introduced as the minimum skew information achievable on a single local measurement on the subsystem $A$ and it has been obtained as
@@
MIN as a measure of nonlocality, a unique property of quantum mechanics, is a description of the maximum global effect caused by locally invariant measurements LuFu-PRL-2011 . Originally, MIN was defined in terms of the Hilbert-Schmidt norm $||X||_{2}=\sqrt{\text{tr}(X^{\dagger}X)}$ but, then, it was argued that this definition suffers the ancilla and noncontractivity problems. Thus, it was defined based on the trace norm HuFa-NJP-2015 , Wigner-Yanase skew information Lietal-EPL-2016 and fidelity MuSa-PLA-2017 . Trace-norm based MIN avoids problems was mentioned above and is defined via
@@
Here, the maximum is taken over the complete set of local projective measurements $\Pi^{(A)}=\{\Pi_{i}^{(A)}\}$, which do not disturb the reduced state $\rho^{(A)}=\text{tr}_{B}(\rho)$ i.e., $\sum_{i}\Pi_{i}^{(A)}\rho^{(A)}\Pi_{i}^{(A)}=\rho^{(A)}$. Trace norm is invariant under local unitary transformation CiTuGi-NJP-2014 . By using proper local unitary transformations, one transforms the X state (6) into
@@
which is still X-shaped but now with real elements EsKhMaDa-OQE-2022 . For such type of X-shaped states, it has been shown that the trace-norm MIN (24) recasts QiLi-QIP-2016
@@
Another measure of quantum correlations is correlated coherence Taetal-PRA-2016 ; Huetal-PR-2018 which is defied as
@@
where $\rho^{(A)}=\text{tr}_{B}(\rho)$ and $\rho^{(B)}=\text{tr}_{A}(\rho)$ are reduced states describing respectively subsystems $A$ and $B$; and $C(\rho)$ is the “quantum coherence” BaPl-PRL-2014 of the state $\rho$ quantified by e.g., $\ell_{1}-$norm which is the sum of the magnitudes of all off-diagonal elements. The key idea in this definition is to split the total coherence contained in the state $\rho$ into the local ones, stored locally in the subsystems, and the nonlocal one stored only in the correlated states. Connection of this notion to quantum discord and entanglement has been argued Taetal-PRA-2016 . For the X state (6), the reduced states are diagonal,

### v2

# Quantum correlations under environmental decoherence

@@ I Introduction
Quantum states are fundamentally different from classical states. Any measurements made on one part of a bipartite state will make the results uncertain. The integral nature of quantum states is the reason for this randomness, not the measuring device. Only states, which can be described by a classical probability distribution, are invariant under such local measurements. Perhaps the most striking feature of quantum mechanics is the quantum entanglement (QE). It is at the heart of quantum information processing, being a resource for doing tasks like quantum computing, quantum teleportation and quantum key distribution [1]. Nonetheless, in the early age of quantum information science, some researchers questioned entanglement as the source of power for quantum computation [2]. QE is not the only correlation which exists between different parts of a multi-party quantum system. Although pure states can be either uncorrelated or just entangled [3], it has been proven that certain quantum separable mixed states exhibit quantum correlations, the so-called quantum discord [4, 5]. In addition to the significant role of quantum correlations in efficient quantum communication and computational tasks, they are also important in figuring out cooperative phenomena in multi-party quantum systems. Among various beyond entanglement quantum correlations have already been identified, one can mention local quantum uncertainty (LQU) [6] and measurement-induced nonlocality (MIN) [7, 8]. While the former is introduced as the minimum skew information [9] achievable on a single local measurement, the latter is interpreted as the maximum global effect caused by locally invariant measurements. To see some applications of quantumness beyond entanglement see [3].
@@
Beside quantum correlations, quantum coherence [10], being cornerstone of quantum computation and quantum communication theory, has a fundamental and practical significance for developing the quantum mechanics [11]. Contrary to quantum correlations, being defined for systems with at least two parties, quantum coherence measures can also be defined for single systems. However, both arise from quantum superposition principle, thus being interrelated. In this regard, the link between quantum coherence and entanglement have already been identified [12, 13]. Furthermore, the concept of “correlated coherence”(CC) [14] has been introduced and its connection to quantum discord and entanglement has been argued. The idea is splitting the coherence contained in the global state into two portions, those stored locally in the marginal (reduced) states and that stored non-locally only in the correlated state, and then subtracting local portions from the correlated one. For X states, we’ll find that MIN equals CC.
@@
Since a quantum system cannot be completely isolated from its environment, its influence on the system’s dynamics must be taken into account. Two approaches have been developed to study dynamics of open quantum systems: the quantum operations formalism and the approach of master equations, which each one has its own place [1]. The latter strategy employs a differential equation, which, in addition to the conventional unitary term for an isolated system, includes a non-unitary term responsible for interaction with the surrounding environment. In the decoherence process, one of the most striking results is the complete loss of entanglement in a finite time, the so-called “entanglement sudden death” (ESD) implying complete termination of entanglement after a finite time interval without a long-lasting smooth decrease [15, 16, 17]. Mathematically, ESD is based on the cutoff in the definition of concurrence, but it is difficult to find a physical interpretation. However, its connection to energy transfer between the system and its surroundings has been discussed [18]. On contrary, the environment-induced [19] or decoherence-induced entanglement [20] has been worked out. Here, environment plays a constructive role, in contrast to its customary destructive role, inducing QE. Furthermore, entanglement re-birth or revival has also been reported. As quantum correlations and particularly quantum entanglement form the basis of quantum technologies, a significant concern pertains to the prevention or management of disentanglement resulting from decoherence in realistic physical systems. Different approaches have been proposed for retarding or avoiding ESD [21, 22, 23, 24].
@@
In this work, we examine a system of two localized coupled qubits that are in interaction with an external magnetic field. The interaction with the environment is introduced by a non-unitary term in the Lindblad equation. We are going to solve this equation analytically first for a number of initial states in the form X and then for a general X-shaped state; and study the evolution of different quantum correlations for different emission rates as well as different interaction intensities. It is imperative to determine whether quantum correlations beyond entanglement experience sudden death or not. Furthermore, it is important to conduct an investigation into the phenomenon of correlation induced by the environment. After all this, it will be time to consider the effect of temperature on quantum correlations. How does this parameter affect ESD?
It is noteworthy that X states have been extensively analyzed in the literature from different aspects. Analytical expressions for diverse measures of quantum correlations have been derived for general X states [25, 26, 27, 28] and their dynamical behavior has been examined. For instance, see [29, 30, 31, 32, 33, 34]. Furthermore, it is worth mentioning that investigation of the effect of decoherence on quantum correlations through other approaches like intrinsic decoherence in the Milburn framework [35, 36] or quantum channels [1] had been the subject of many studies [37, 38, 39, 40, 41, 42, 43].
@@ II Some Quantum Correlations; entanglement and beyond
In this section we briefly review QE, LQU and MIN. At the end, we consider CC and relate it to MIN. Different measures have been introduced for quantification of QE. One of the most famous measures of entanglement is the concurrence [44, 45] which for a two-qubit state $\rho$ is defined as
@@
$\lambda_{i}$s being the eigenvalues of the Hermitian matrix $\sqrt{\sqrt{\rho}\tilde{\rho}\sqrt{\rho}}$ in non-increasing order. Here, $\tilde{\rho}$ is the spin-flipped Hermitian density matrix given by $(\sigma_{y}\otimes\sigma_{y})\rho^{*}(\sigma_{y}\otimes\sigma_{y})$; $\sigma_{y}$ being the second Pauli matrix and $\rho^{*}$ the complex conjugate of $\rho$ which due to the Hermiticity of $\rho$ equals its transpose, $\rho^{*}=\rho^{T}$. Two matrices $S=\sqrt{\rho}\tilde{\rho}\sqrt{\rho}$ and $R=\rho\tilde{\rho}$ are connected through the invertible matrix $\sqrt{\rho}$ i.e., $R=(\sqrt{\rho})S(\sqrt{\rho})^{-1}$. Thus, they are similar, having the same spectrum. In this way, one sees $\lambda_{i}$s are the square roots of the eigenvalues of $S$ and thus the non-Hermitian matrix $R$. Note that since $\sigma_{y}\otimes\sigma_{y}$ is its own inverse, $(\sigma_{y}\otimes\sigma_{y})^{-1}=\sigma_{y}\otimes\sigma_{y}$ then $\tilde{\rho}$ is similar to $\rho^{*}$ which its eigenvalues are just those of $\rho$. Therefore, both $\rho^{*}$ and its similar matrix $\tilde{\rho}$ are non-negative. Now, since both $\rho$ and $\tilde{\rho}$ are non-negative, eigenvalues of their product $R=\rho\tilde{\rho}$ are non-negative [46] [^1] See page 256.. Concurrence, defined in this way, becomes zero for separated states but takes the value one for maximally entangled states in $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ systems [47]. It has been argued that the concurrence is related to the entanglement of formation, which itself is an entropic entanglement monotone [44, 45]. In the basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$, a X-state has the form
@@
coming from properties of the density operator i.e., unit trace, Hermiticity and positivity. The concurrence (1) for the X-state (2) is then [48]
$\mathcal{C}$ $=$ $\max\{0,\mathcal{C}_{1},\mathcal{C}_{2}\}$ (4)
where,
$\mathcal{C}_{1}=2(|\rho_{14}|-\sqrt{\rho_{22}\rho_{33}})$ (5)
$\mathcal{C}_{2}=2(|\rho_{32}|-\sqrt{\rho_{11}\rho_{44}})$ (6)
As these equations show, if and only if either $\rho_{22}\rho_{33}<|\rho_{14}|^{2}$ or $\rho_{11}\rho_{44}<|\rho_{23}|^{2}$ then the state (2) is entangled. Both conditions cannot be fulfilled simultaneously [49]. Our model which will be considered in section III can be regarded as two effective two-level atoms that are subject to decoherence. As such, $\mathcal{C}_{1}$ provides a measure of an entanglement produced by the two-photon coherence that is between the ground state, with no photons, and the two-photon excited state ($|11\rangle\leftrightarrow|00\rangle$), whereas $\mathcal{C}_{2}$ provides a measure of an entanglement produced by the one-photon coherence corresponding to the coherence that is present between two different one-photon states ($|10\rangle\leftrightarrow|01\rangle$). If an entanglement is generated among the one-photon states, the magnitude of the entanglement is restricted by the populations of the no-photon and two-photon states, and conversely, if an entanglement is created among the two-photon states, the magnitude of the entanglement is restricted by the population of the one-photon states $\rho_{22}$ and $\rho_{33}$ [50].
The basis of collective states of the system, or the Dicke states, which are defined as
$|e\rangle=|00\rangle$ (7)
$|g\rangle=|11\rangle$ (8)
$|s\rangle=(|01\rangle+|10\rangle)/\sqrt{2}$ (9)
$|a\rangle=(|01\rangle-|10\rangle)/\sqrt{2}$ (10)
is a particularly useful choice for working with. Here, the excited and ground states $|e\rangle$ and $|g\rangle$ are separable, while the symmetric and antisymmetric states $|s\rangle$ and $|a\rangle$ are maximally entangled. The X-state (2) in this new basis is then expressed as,
$\varrho$ $=$ $\left(\begin{array}[]{cccc}\rho_{11}&\rho_{14}&0&0\\ \rho_{41}&\rho_{44}&0&0\\ 0&0&\frac{1}{2}(\rho_{22}+\rho_{33})+\text{Re}\{\rho_{32}\}&\frac{1}{2}(\rho_{22}-\rho_{33})+i~{}\text{Im}\{\rho_{32}\}\\ 0&0&\frac{1}{2}(\rho_{22}-\rho_{33})-i~{}\text{Im}\{\rho_{32}\}&\frac{1}{2}(\rho_{22}+\rho_{33})-\text{Re}\{\rho_{32}\}\\ \end{array}\right)\equiv\left(\begin{array}[]{cccc}\varrho_{ee}&\varrho_{eg}&0&0\\ \varrho_{ge}&\varrho_{gg}&0&0\\ 0&0&\varrho_{ss}&\varrho_{sa}\\ 0&0&\varrho_{as}&\varrho_{aa}\\ \end{array}\right)$ (19)
which has the block-diagonal form. Note that the diagonal elements are interpreted as populations, while the off-diagonal ones represent coherences [66]. In terms of the density matrix in the Dicke basis, $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ of Eq. (4) is given by [66, 67]
$\mathcal{C}_{1}$ $=$ $2\left\{|\varrho_{eg}|-\left[\left(\frac{\varrho_{ss}+\varrho_{aa}}{2}\right)^{2}-(\text{Re}\{\varrho_{sa}\})^{2}\right]^{1/2}\right\}$ (20)
$\mathcal{C}_{2}$ $=$ $2\left\{\left[\left(\frac{\varrho_{ss}-\varrho_{aa}}{2}\right)^{2}+(\text{Im}\{\varrho_{sa}\})^{2}\right]^{1/2}-\sqrt{\varrho_{ee}\varrho_{gg}}\right\}$ (21)
As another natural and widely used entanglement measure, negativity for a bipartite state $\rho$ is defined through the trace norm of the partially transposed density matrix [51],
@@
where $\{|i\rangle\}$ and $\{|k\rangle\}$ are given local orthonormal basis for subsystems $A$ and $B$ respectively. It is notable that the spectrum of the partial transposition of the density matrix is independent both of the choice of the local basis and also whether the partial transposition is performed over the party $A$ or $B$ [52]. One can easily check that $\rho^{T_{B}}$ is Hermitian
@@
where in the third equality we have used Hermiticity of the density operator $\rho$. Now, from Hermiticity of $\rho^{T_{B}}$ and the fact that trace is independent of basis, one has that
@@
where in the third equality we have used the fact that for entangled two qubit states the partiality transposed state $\rho^{T_{B}}$ has only one negative eigenvalue [53] which has been denoted by $\lambda_{\min}$ in Eq. (25) as it is the minimum eigenvalue. Some authors have defined negativity of two-qubit states as
@@
which for maximally entangled states where the spectrum of $\lambda$ is $\{-\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2}\}$, takes the largest value 1 [54, 55]. For a mixed state of two qubits, the subject of our study, it has been proved that the negativity $\mathcal{N}^{\prime}$ of a state can never exceed its concurrence and is always greater than $\sqrt{(1-\mathcal{C})^{2}+\mathcal{C}^{2}}-(1-\mathcal{C})$; $\mathcal{C}$ being the concurrence of the state [55]. Now, from Eqs. (25) and (26) we have that $\mathcal{N}=\mathcal{N}^{\prime}/2$ and thus from the results of [55] we have that
@@
The logarithmic negativity (log-negativity) as an appropriate quantifier of QE in a bipartite quantum system is defined as [51, 56]
@@
Now we consider quantum correlations beyond QE. Based of the notion of skew information, authors of [6] have proposed a discord-like scheme for characterizing non-classical correlation via LQU. In this way, LQU has been introduced as the minimum skew information achievable on a single local measurement on the subsystem $A$ and it has been obtained as
@@
MIN as a measure of nonlocality, a unique property of quantum mechanics, is a description of the maximum global effect caused by locally invariant measurements [7]. Originally, MIN was defined in terms of the Hilbert-Schmidt norm $||X||_{2}=\sqrt{\text{tr}(X^{\dagger}X)}$ but, then, it was argued that this definition suffers the ancilla and noncontractivity problems. Thus, it was defined based on the trace norm [8], Wigner-Yanase skew information [57] and fidelity [58]. Trace-norm based MIN avoids problems was mentioned above and is defined via
@@
Here, the maximum is taken over the complete set of local projective measurements $\Pi^{(A)}=\{\Pi_{i}^{(A)}\}$, which do not disturb the reduced state $\rho^{(A)}=\text{tr}_{B}(\rho)$ i.e., $\sum_{i}\Pi_{i}^{(A)}\rho^{(A)}\Pi_{i}^{(A)}=\rho^{(A)}$. Trace norm is invariant under local unitary transformation [59]. By using proper local unitary transformations, one transforms the X state (2) into
@@
which is still X-shaped but now with real elements [60]. For such type of X-shaped states, it has been shown that the trace-norm MIN (34) recasts [61]
@@
Another measure of quantum correlations is correlated coherence [14, 11] which is defied as
@@
where $\rho^{(A)}=\text{tr}_{B}(\rho)$ and $\rho^{(B)}=\text{tr}_{A}(\rho)$ are reduced states describing respectively subsystems $A$ and $B$; and $C(\rho)$ is the “quantum coherence” [10] of the state $\rho$ quantified by e.g., $\ell_{1}-$norm which is the sum of the magnitudes of all off-diagonal elements. The key idea in this definition is to split the total coherence contained in the state $\rho$ into the local ones, stored locally in the subsystems, and the nonlocal one stored only in the correlated states. Connection of this notion to quantum discord and entanglement has been argued [14]. For the X state (2), the reduced states are diagonal,